OpenStudy (anonymous):

Express the given quantity as a single logarithm: ln(1+x^5)+1/2lnx-lncosx

OpenStudy (tkhunny):

Have you considered using the logarithm rules? ln(a) + ln(b) = ln(ab) aln(b) = ln(b^a) ln(a) - ln(b) = ln(a/b) That's all you need.

OpenStudy (anonymous):

Yes, but i still don't understand how to apply the rules to this question.

OpenStudy (tkhunny):

Okay, start with the middle term, just for fun. (1/2)ln(x) Look really hard at the middle rule. Can you see how it applies?

OpenStudy (anonymous):

(1/2)ln(x) = to lnx^(1/2)?

OpenStudy (tkhunny):

Perfect. No you have: ln(1+x^5)+lnx^½-lncosx Okay, look at the first two terms. How can you apply the first property?

OpenStudy (anonymous):

ln+lnx^5+lnx^1/2-lncosx?

OpenStudy (anonymous):

oh wait...lnx^5+lnx^1/2-lncosx?

OpenStudy (tkhunny):

No good. You cannot break up addition inside a logarithm. There is no property like that. ln(1+x^5)+ln(x^½)-lncosx Using the first rule: ln[(1+x^5)*(x^½)]-lncosx Now, use the last rule to finish up.

OpenStudy (anonymous):

ln(1+x^5)*(lnx^1/2)/-lncosx?

myininaya (myininaya):

He wants you to use ln(a)-ln(b)=ln(a/b)

OpenStudy (tkhunny):

ln(a) + ln(b) - ln(c) = ln(a*b) - ln(c) = ln((a*b)/c)